OpenStudy (anonymous):

how to graph a line x-y= -4

OpenStudy (anonymous):

i would change it to the form y=mx+b

OpenStudy (anonymous):

ok can you walk me through this one and then me see if i understand it

OpenStudy (anonymous):

subtract x from both sides, divide by negitive one to get you a postive y

OpenStudy (saifoo.khan):

x - y = -4 When u have to plot the "x-intercept", put y = 0 When u have to plot the "y-intercept", put x = 0 So, the x intercept, x - y = -4 x - 0 = -4 x = -4 Now the y intercept, x - y = -4 0 - y = -4 y = 4 So the line will cut the x axis at -4 and y at 4.

OpenStudy (saifoo.khan):



OpenStudy (anonymous):

3x+2y=0

OpenStudy (anonymous):

i don't understand how to do it

OpenStudy (saifoo.khan):

Did u get the first thing?

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

Why?

OpenStudy (anonymous):

i don't know. i don't get how you got -4 and 4

OpenStudy (saifoo.khan):

I did step by step.. read it again.. and try to understand/..

OpenStudy (anonymous):

ok. let me see if i get this right: 3x+2y=0 x crosses at 3 y crosses at 2

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

Nope.. not like that.. This is a bit different case..in the second one the line passes through the Origin.

OpenStudy (saifoo.khan):

First try to understand the first one..

OpenStudy (anonymous):

i get it. can you explain the second one?

OpenStudy (saifoo.khan):

Alright.. so in this one, u have do a bit differently, Coz when u are going to put the x zero to find the y intercept.. u will get zero.. and same as with the x intercept.. So this thing proves that the line will cut the axes at (0,0) At that point..: